OpenStudy (anonymous):

What i sthe range of the quadratic function F(x)= -5x^2+30x+4 A.) Y<49 B.) Y<3 C.) x>49 D.) Y>49

OpenStudy (anonymous):

range is always something with y, so C is wrong. The range depends on the minimum or maximum of the function. To determine the maximum or minimum of a function, one uses the derivative and sets it to 0. f(x)=-5x²+30x+4 f'(x)=-10x+30 f'(x)=0 -10x+30=0 30=10x x=3 Let's fill in x=3 in f(x): f(3)=-5*3²+30*3+4=-5*9+94=94-45=49 Now, does f(x) have a maximum or a minimum?

OpenStudy (anonymous):

si it has to be B

OpenStudy (anonymous):

No, f(x)=y and y=f(3)=49, so the top can't be on y=3.

OpenStudy (anonymous):

ohh okay C

OpenStudy (anonymous):

Are you just guessing now? I already stated C was wrong.

OpenStudy (anonymous):

ohhh i could swear if > is greater then right

OpenStudy (anonymous):

you there i just relize that this is less> and this is greater <so the answer is D